OpenStudy (anonymous):

a=4,b=−8,c=1 x1=1+1/2 √3 x2= 1-1/2 √3 Could anyone PLEASE walk me through this?

jimthompson5910 (jim_thompson5910):

you want to use the quadratic formula to solve 4x^2 - 8x + 1 = 0 right?

OpenStudy (anonymous):

Right

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

Plug in the given values a = 4 b = -8 c = 1 to get \[\Large x = \frac{-(-8)\pm\sqrt{(-8)^2-4(4)(1)}}{2(4)}\] Make sense so far?

OpenStudy (anonymous):

I am following

jimthompson5910 (jim_thompson5910):

ok now evaluate (-8)^2-4(4)(1) to get 48 and simplify -(-8) to get 8 so we get this next \[\Large x = \frac{8\pm\sqrt{48}}{8}\]

jimthompson5910 (jim_thompson5910):

from here, break up the plus/minus to get \[\Large x = \frac{8+\sqrt{48}}{8} \ \text{or} \ x = \frac{8-\sqrt{48}}{8}\]

jimthompson5910 (jim_thompson5910):

and simplify \[\Large \sqrt{48}\] to get \[\Large 4\sqrt{3}\] so after that, the next step is \[\Large x = \frac{8+4\sqrt{3}}{8} \ \text{or} \ x = \frac{8-4\sqrt{3}}{8}\] do you see what to do from here?

OpenStudy (anonymous):

divide by 8?

jimthompson5910 (jim_thompson5910):

yes divide each part by 8 or break up the fractions and reduce \[\Large x = \frac{8}{8}+\frac{4\sqrt{3}}{8} \ \text{or} \ x = \frac{8}{8}-\frac{4\sqrt{3}}{8}\] \[\Large x = 1+\frac{1}{2}\sqrt{3} \ \text{or} \ x = 1-\frac{1}{2}\sqrt{3}\]

OpenStudy (anonymous):

Thank you so much easier with step by step help

jimthompson5910 (jim_thompson5910):

yw